Question 1 (25 pts)

Find the running time complexity for the following code fragments. Express your answers using either the Big-O or Big-Θ notations, and the tightest bound possible. Justify your answers.

for(int count O , i -0; i < n* n; i++) for(int i0 ; j <i; j++) count++
Answer #1
First loop is running for n*n times.
Second for loop is running for i times.
Max value of i is n*n.

So, Total time complexity = O(n*n*n*n) = O(n^4)